The following tables show a small firm’s long-run average cost of manufacturing a good at two different plants:

Plant 1:

Quantity: 1 2 3 4 5 6 7 8 9

T.C.: 50 106 164 224 287 355 430 520 618

A.C. ? ? ? ? ? ? ? ? ?

M.C. ? ? ? ? ? ? ? ? ?

Plant 2:

Quantity: 1 2 3 4 5 6 7 8 9

T.C.: 20 52 90 130 175 227 285 345 407

A.C. ? ? ? ? ? ? ? ? ?

M.C. ? ? ? ? ? ? ? ? ?

T.C.=Total Cost

A.C.=Average Cost

M.C.=Marginal Cost

a) Complete the third and fourth columns of each table.

b) Suppose the price of the good is $60. How much should the firm produce in each plant in order to maximize the firm’s profit? Find the firm’s profit.



1
Expert's answer
2021-12-07T10:23:24-0500


Average cost ="\frac{Total cost}{Number of units }"

Marginal cost = "\frac{Change in cost}{Change in quantity}"

(b)

Maximum Profit occurs at MR=MC

MC= 60

For Plant 1 Maximum profit is achieved by producing 6 units while for Plant 2 Maximum profit is achieved by producing 8 units

Profit= TR-TC

For Plant 1; The profit is 16

For Plant 2; The profit is 135